INFRARED SPECTROSCOPY OF THE MIXED N_{2}O-CO_{2} DIMER

\underline{M. AFSHARI}, M. DEHGHANY, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, 2500 University Drive NW, Calgary, Alberta T2N 1N4, Canada; A.R.W. MCKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

High-resolusion infrared spectra of the weakly bound N_{2}O-CO_{2} complex are studied using a tunable diode laser to probe a pulsed supersonic slit jet. The previously known N_{2}O-CO_{2} dimer band in the region of the ν_{3} CO_{2} asymmetric strtch (∼ 2350 cm^{−1}) is remeasured and analyzed in improved detail and two new bands in the regions of the N_{2}O ν_{1} (∼ 2230 cm^{−1}) and ν_{3} (∼ 1280 cm^{−1}) stretching fundamentals are observed and assigned to N_{2}O-CO_{2}. The ground state rotational constants for all three bands are A = 0.29498 cm^{−1}, B = 0.05801 cm^{−1} and C = 0.04837 cm^{−1}. We also observe another band with c-type rotational structure at about 2251.5 cm^{−1} which is assigned as a combination of the intramolecular N_{2}O ν_{1} stretching vibration and the intermolecular out-of-plane torsional vibration. The resulting torsional frequency for the N_{2}O-CO_{2} dimer is about 25.7 cm^{−1}.